A force of 25 N accelerates a body by 5 m/s2. What is the mass of the body?
1
Expert's answer
2012-11-13T11:51:07-0500
According to the second Newton's law the mass of the body is

m = F/a = 25[N]/5[m/s²] = 5[kg].